Logiweb dictionary of peano commutativity

Up Help

0 0 peano commutativity

1 0 Weakening 2 0 ConditionedMP′ 3 0 PermuteAntecedents 4 0 ImplyTransitivity 5 0 EqualReflexivity 6 0 EqualSymmetry 7 0 EqualTransitivity 8 0 Mendelson3.2(d) 9 0 SubstitutionMacro 10 0 Mendelson3.2(f)Base 11 0 Mendelson3.2(f)Indu 12 0 Mendelson3.2(f) 13 0 Mendelson3.2(g)Base 14 0 Mendelson3.2(g)Indu 15 0 Mendelson3.2(g) 16 0 Mendelson3.2(g)Switch 17 0 AddOne 18 0 Mendelson3.2(d)Rule 19 0 Mendelson3.2(d)CondRule 20 0 EqualTransitivityRule 21 0 EqualTransitivityCondRule 22 0 PlusCommutativityBase 23 0 PlusCommutativityIndu 24 0 PlusCommutativity 25 0 DoubleMP′ 26 0 InductionMacro 27 0 EqualSymmetryRule 28 0 DoubleConditionedMP′ The pyk compiler, version 0.grue.20050603 by Klaus Grue GRD-2005-06-29.UTC:11:26:23.740136 = MJD-53550.TAI:11:26:55.740136 = LGT-4626761215740136e-6